HUBBLE SPACE TELESCOPE DAILY REPORT #3073 PERIOD COVERED: 0000Z (UTC) 03/14/02 - 0000Z (UTC) 03/15/02 Daily Status Report as of 074/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/MA1/MA2 8959 (MAMA Dark vs. Temperature) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to measure the dark rate for both MAMA detectors as a function of instantaneous temperature and recent thermal history. These will be used to make measurements at the beginning and end of three consecutive SAA-free orbit blocks. The proposal completed with no reported problems. 1.2 Completed STIS/CCD/MA1/MA2 8961 (End of BEA Test) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to assess whether STIS sensitivity is significantly below pre-SM3 performance by using two G140L spectra of flux standards. Significantly degraded sensitivity will trigger additional observations of flux standards to validate the initial result and/or track subsequent decontamination. All standards used in this program will have recent preparatory observations obtained in the same configuration {8956}. Measurement accuracy will be limited by global sensitivity fluctuations {less than 5%} rather than photon counting statistics. There was no reported anomalous activity. 1.3 Completed ACS/WFC 9006 (CCD Temperature Set Point Determination) The Advanced Camera for Surveys (WFC) was used to adjust the CCD temperature to nominal operating level. The proposal completed nominally. 1.4 Completed Two Sets of STIS/CCD 8963 (CCD Bias) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD read noise from multiple bias frames obtained in common science modes: amplifier D, gains 1 and 4 electrons/DN, and binning factors of 1x1, 1x2, 2x1, and 2x2. The read noise measurements before and after NCC turn-on are then compared, and standard calibration reference files for use in pipeline reduction of SMOV3B data are constructed. The proposal had no reported problems. 1.5 Completed Three Sets of FGS/1 9581 (A BEA Science Program: A High Angular Resolution Survey of Some of the Galaxy's Most Massive Stars) Fine Guidance Sensor #1 was used to make optimal use of otherwise unused telescope time while HST's pointing is restricted to the BEA cone during the first 12 days of SMOV3B. The launch date of Feb 29 placed two of the Galaxy's most active star forming regions, Trumplers 14 and 16 in the Carina Nebula, within the BEA for the entire 12 day period. FGS1r was used to conduct a high angular resolution survey of the most massive and luminous stars in these two magnificent clusters to establish the incidence of binarity at the high mass end of the HR diagram. This survey will close the detection gap between spectroscopic and speckle surveys in which duplicity can be concealed and will dramatically improve our knowledge of the IMF at the high mass end of the main sequence as well as the luminosity function {LF} of star forming regions which produce such high mass stars. Finally, this survey will identify targets for an approved SIM Key Science project to dynamically determine O-star masses. There was no anomalous activity. 1.6 Completed STIS/CCD 8964 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain dark frames twice daily, as in a corresponding calibration program. Standard calibration reference darks are constructed and hot pixel lists are updated for use in pipeline reduction of SMOV3B data. All observations completed without incident. 1.7 Completed STIS/CCD 8958 (Image Quality) The Space Telescope Imaging Spectrograph (CCD) was used to check image quality using two tests. In the first test, an ACQ/PEAK was used to dither the 0.1x0.09 aperture across a point source to sample the PSF. Light was dispersed entering the spectrograph onto the CCD using the G230LB grating. Relative counts at different wavelengths will give a measure of PSF shape at different wavelengths. In the second test, a point source is imaged onto the CCD through the F28x50OII filter. The resulting PSF is then compared to the PSF shape prior to SM3B. No problems were noted. 1.8 Completed Five Sets of ASC/HRC 9005 (CCD Functional) The Advanced Camera for Surveys (HRC) was used to measure the baseline performance and commandability of the CCD subsystems using exposures of bias frames, darks and flat fields with tungsten lamp. The on-board baseline compression algorithm for WFC data are measured. This procedure is carried out purely with internal targets and observations do not need to be contiguous. No anomalous activity occurred. 1.9 Completed HST 8968 (Vehicle Disturbance Test) The Vehicle Disturbance Test {VDT} was used to characterize uncompensated environmental disturbances acting upon the HST during normal operation. The VDT is a passive test {not a forced-response test} used to obtain signatures for both externally induced {e.g. SA-3} and internally induced {e.g. NCC and ACS mechanisms} disturbances for comparison with past VDT results. The disturbances observed will be used as the nominal on-orbit disturbances in pointing control simulations until the next VDT is run. The test occurs after release, and most of the VDT can be run during the BEA period. The -V1 sunpoint portion of the VDT occurs after the BEA period is complete. There were no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 6 Successful: 6 Scheduled Re-acquisitions: 0 Successful: 0 2.2 FHST Updates: Scheduled: 12 Successful: 12 2.3 Operations Notes: QDV10GT1 (10 Second GG Torque 2) flagged out-of-limits at 073/131620Z. The menmonic was back in bounds at 073/131832Z. HSTAR 8552 was written. Operations requests were used to initiate: the ACR setup for neon leak high voltage protection at 073/1340Z the NSSC-1 high voltage protect table modification at 073/1642Z the Flight Software 5.8.3 installation at 073/1835Z this version updates the DT temperature limit and mitigates the potential SEU-induced spurious command interrupt the ASCS/NCS quick update for M8FSW83 at 073/2215Z The engineering status buffer limit for SESBSLD was updated to '4' at 073/1626Z, using ROP DF-18A. The Vehicle Disturbance Test (VDT) #2 (see 1.9 above) executed nominally between 073/225000Z and 074/065000Z. VDT #2 was a 5-orbit test at +V3 sunpoint designed to establish HST baseline jitter characteristics prior to NCS startup. This test will be repeated next week as VDT #4 after NCS cooldown to observe differences attributable to disturbances from the NICMOS Cooling System. Similar to VDT #1, the largest disturbances observed during VDT #2 were reaction wheel zero-speed crossings with amplitudes not exceeding 50 mas. Additional disturbances were observed during mid-day and mid-night intervals in the HST V2 rotation axis having an amplitude of ~25 mas lasting 5-minutes or less that were not correlated with HGA motion, RWA zero-speed crossings, or EOD/EON terminators Spectral post-processing of the test data will be performed to identify the source of these events. The control law gains (maneuver gains) used during the VDT amplify the HST jitter response by a factor of 10 above the nominal control law (acquisition) gains, so these unidentified disturbances will have an amplitude of ~2.5 mas when using the nominal post-SM3B control law. VDT #3 Contingency and the Nominal VDT #3 will begin on Saturday evening at 11:30 PM local time and Sunday morning at 8:00 AM, respectively. The purpose of VDT #3 is to capture the NCC startup surging disturbances shortly after NCS activation. In support of the VDT #2, there were ten LGBU data takes per the dictates of ROP PO-5. The flight operations team returned to the GSFC MOR to support SMOV activities throughout the upcoming weekend. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW